Please help! It would be greatly appreciated!

Comprehensive Accounting Cycle Review 11-2 Karen Noonan opened Larkspur Inc. on February 1, 2017. During February, the following transactions were completed: Feb. 1 Issued 4,000 shares of Larkspur common stock for $13,260. Each share has a $1.50 par. Borrowed $8,000 on a 2-year, 6% note payable Paid $9,200 to purchase used floor and window cleaning equipment from a company going out of business ($5,120 was for the floor equipment and $4,080 for the window equipment) 1 1 Paid $220 for February Internet and phone services 3 Purchased cleaning supplies for $1,000 on account. 4. Hired 4 employees. Each will be paid $490 per 5-day work week (Monday- Friday). Employees will begin working Monday, February 9 Obtained insurance coverage for $11,040 per year. Coverage runs from February 1, 2017, through January 31, 2018. Karen paid $2,760 cash for the first quarter of coverage Discussions with the insurance agent indicated that providing outside window cleaning services would cost too much to insure. Karen sold the window cleaning equipment for $4,030 cash. Billed customers $3,980 for cleaning services performed through February 13, 2017. Received $540 from a customer for 4 weeks of cleaning services to begin February 21, 2017. Paid $310 on amount owed on cleaning supplies Paid $3 per share to buy 300 shares of Larkspur common stock from a shareholder who disagreed with management goals 5 16 17 18 20 23 24 25 27 28 The shares will be held as treasury shares Billed customers $4,390 for cleaning services performed through February 20 Paid cash for employees wages for 2 weeks (February 9-13 and 16-20) Collected $2,550 cash from customers billed on February 16 Paid $220 for Internet and phone services for March Declared and paid a cash dividend of $0.15 per share
Answer #1

1)Account receivbale(db)3880
service revenue(cr)3880
2)Allowance for doubtful account(db)200
account receivable(Cr)200
total receivables=3980+4390-2550=5820
3)bad debt expense(dB) (5820*3%)+200=374.6
Allowance for doubtful account(cr)374.6
4)depreciation expense(db) (5120-800)/(4*12)=90
acc depreciation -equipment(cr)90
5)insurance expnse(db)11040/12=920
prepaid isnurance(db)920
6)supplies expense(db)1000-410=590
supplies(cr)590
7)unearnd service revenue(db)540/4=135
service revenue(Cr)135
8)salaries and wag expense(db)490*4=1960
salaries and wages payable(Cr)1960
9)interest expense(db)8000*6%*(1/12)=40
interest payable(cr)40
10)income tax expense(db)
income tax payabel(Cr)